# Gentle spaghettification in effective LQG dust collapse

### v1

@@ V Tidal forces during effective stellar collapse
As pointed out before, effective inhomogeneous dust collapse does not end in the classical central singularity, which is replaced by a quantum gravitational bounce. In classical Einstein theory, angular tidal forces close to shell-crossing singularities do not diverge, while despite radial tidal forces being divergent, radial deviation does not go to zero at SCS, nor does it diverge Szekeres:1995gy . This makes such singularities less pathological than the central one. Here, we show that similar results hold also at the effective level.
To prove this, we compute explicitly (16), (17), using the solution (3) of the effective equation (2). Let’s start from the equation for angular deviation. After some computation, we find
@@
Firstly, we notice that in the classical case $\Delta\rightarrow 0$, the angular tidal forces negatively diverges for $t\rightarrow\alpha(R)$ (classical central singularity). In contrast, here it remains always bounded, reaches its maximum at $t=\alpha(R)$ (the bouncing point), and two zeroes, for $t=\alpha\pm 2\sqrt{\Delta}/\sqrt{3}$.
@@
Figure 1: Angular tidal force for $\alpha=1$, in the classical (dark blue curve) and effective collapse (red curve). In the classical case the function diverges at the crush singularity $t=\alpha=1$, while remains bounded at the effective level.
Fig.1 shows the comparison between the classical and effective cases. Notice that angular tidal forces do not diverge at shell-crossing singularity, which arises within a time $\alpha(R)<t\leq\alpha(R)+2\sqrt{\Delta}/3$.
@@
Let’s now examine the radial tidal force. By explicitly writing equation (16), after a lengthy but straightforward computation, we find
@@
Firstly, we notice that divergence can only occur if $\partial_{R}r=0$, provided the numerator of the last term is nonzero. This indicates that the radial tidal force remains finite at the bounce. To study the radial tidal force near and at SCS, we can assume that the first term is negligible
@@
Next, we recall that for the profiles we are considering (decreasing, continuous and of compact support), shell-crossing singularities arise for $\alpha<t\leq\alpha+2\sqrt{\Delta}/3$, which implies $F\Delta<r^{3}\leq 2F\Delta$, from (3). This means that the numerator of (21) is strictly positive, and at shell-crossing singularity the radial tidal force diverges. It is worth noting that such a divergence arises in the deep quantum gravitational regime and is a positive divergence: an infinite stretching radial force arises, conceptually similar to the one occurring classically near the central singularity, producing the spaghettification effect. In contrast, near classical shell-crossing singularities, the radial tidal force is negatively divergent: two points radially separated tend to be crushed together by the tidal force, even though in a weak fashion Szekeres:1995gy , as one can see by setting $\Delta=0$ in (21).
Let us now focus on the (positive) divergent behavior of the radial tidal force near and at the SCS. The condition for SCS can be rewritten as Fazzini:2023ova
$(t-\alpha)^{2}\partial_{R}m-2m(t-\alpha)\partial_{R}\alpha+\frac{4\Delta}{9}\partial_{R}m=0~{}.$ (22)
This is a second-degree equation in $t-\alpha$, with solutions $t_{1}=\alpha(R)+\delta t_{1}(R)$, $t_{2}=\alpha(R)+\delta t_{2}(R)$, with $\delta t_{1}\leq\delta t_{2}$. These solutions are guaranteed to be real for continuous, inhomogeneous, decreasing profiles of compact support, as shown in the previous section. It can be also easily shown that the solutions are distinct Fazzini:2023ova , except for an initial condition that satisfies $\frac{m\partial_{R}\alpha}{\partial_{R}m}=\frac{2\sqrt{\Delta}}{3}$, which is a fine-tuned case we do not consider here. Therefore, we can rewrite the condition for SCS as
@@
This means that close to $t_{SCS}=\alpha+\delta t_{1}$, the function $r^{\prime}$ goes to zero linearly: $r^{\prime}\sim K(t_{SCS},R)(t_{SCS}-t)$ for $t\lesssim t_{SCS}$, with $K(t_{SCS},R)>0$. Here we are only interested in tidal forces before and at the shell-crossing singularity. Beyond this point, the equation of motion in differential form breaks down, and analytic results are not reliable. With this in hand, we can rewrite (16):
$\partial^{2}_{t}\delta x^{1}\sim\frac{B(t_{SCS},R)}{t_{SCS}-t}\delta x^{1}~{},$ (24)
where $B(t_{SCS},R)$ is given by $\partial^{2}_{t}\partial_{R}r$ evaluated at shell-crossing singularity. This approximation holds only close to the SCS ($t\lesssim\alpha(R)+\delta t_{1}(R)$). Equation (24) can be solved analytically. Let us define $\delta t\equiv t_{SCS}-t>0$. The solution is
$\delta x^{1}(\delta{t})\sim$ $c_{1}\sqrt{\delta{t}}~{}I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]+$ (25)
$c_{2}\sqrt{\delta{t}}~{}K_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]~{},$ (26)
where $I_{1}$ and $K_{1}$ are modified Bessel functions of the first and second kinds, and $c_{1},c_{2}$ are constants determined by initial conditions. The previous solution holds only for $\delta{t}\rightarrow 0^{+}$. While the first term goes to zero in this limit, the second term approaches a constant non-zero value. Their explicit Taylor expansions are:
@@
Therefore, even though the radial tidal forces diverge at SCS, two neighboring points that radially separated (and thus belong to two different matter layers) do not clash in general when shell-crossing singularity occur. For instance, these points can be thought of as belonging to two shells close to the ones that cross, but with a larger coordinate separation $\delta R$). If these points belong exactly to the crossing shells, they will clearly be crushed, and this condition is expected to be given by $c_{2}=0$ Szekeres:1995gy . This implies that the radial portions of the collapsing star (or even geodesic test particles radially separated) undergo spaghettification, but without being infinitely stretched. We refer to this effect as gentle spaghettification.
This result leads to the conclusion that SCS, which generally arise in effective stellar collapse, are less pathological than classical crush singularities. It also justifies the extension of the spacetime beyond SCS Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf , as done in the classical case through weak solutions. It is important to emphasize that this result is non-trivial. As mentioned before, even though the structure of the deviation equation follows the classical form, the effective solution differs from the classical one. Therefore, the behavior of tidal forces at shell-crossing singularities also differs from the classical behavior in the effective model.
To conclude, for profiles that develop shell-crossing singularities in the pre-bounce phase (where classical Einstein equations hold), the classical analysis Szekeres:1995gy can be directly exported to the effective model.
@@
Shell-crossing singularities are generally unavoidable in effective dust collapse inspired by loop quantum gravity, within the $\bar{\mu}+K$-loop quantization scheme. They therefore play a central role in such effective models, and their geometrical and physical features need to be investigated. In this work, we focused on the analysis of the tidal forces arising from effective dust collapse, with particular attention to the bouncing point and shell-crossing singularities. We found that tidal forces remain bounded at the bounce, in contrast to what happens in the classical case at the central crush singularity, and the same holds for the angular component at the shell-crossing singularity. However, the radial component of the tidal force positively diverges at the shell-crossing singularity.
@@
Nonetheless, when studying the behavior of the radial deviation by solving the Jacobi equation, we find that it remains bounded during the whole dynamics, also when the shell-crossing singularity arises. The positive divergent behavior of the tidal force implies a spaghettification effect, similar to the one occurring near the central singularity in classical stellar collapse. However, the bounded behavior of the radial displacement between nearby points makes such an effect much less dramatic. We refer to this phenomenon as gentle spaghettification, and it allows the dynamics to extend beyond shell-crossing singularities in a meaningful way, through the integral form of the equations of motion. Future work in this direction will aim to investigate tidal forces in the effective gravitational collapse of dust in the non-marginally bound case Cipriani:2024nhx , and of fluids with pressure Cafaro:2024lre , in order to understand, at least at the numerical level, the role of pressure in the tidal forces experienced by test particles.
@@ Acknowledgements.
Appendix A Effective geodesic flow of dust particles
At the effective level the energy-momentum tensor is covariantly conserved
$\nabla^{\mu}T_{\mu\nu}=0~{}.$ (29)
This is a direct consequence of the fact that the matter part of the Einstein equation is not modified by quantum corrections. The energy-momentum tensor for dust reads $T_{\mu\nu}=\rho u_{\mu}u_{\nu}$, where $u^{\mu}$ is the fluid $4$-velocity in its rest frame. By imposing (29) one can easily get $\rho u^{\mu}\nabla_{\mu}u^{\nu}=0$, which for $\rho\neq 0$ implies that dust particles follow geodesics also at the effective level.

### v2

@@ V Tidal forces during effective stellar collapse
As pointed out before, effective inhomogeneous dust collapse does not end in the classical central singularity, which is replaced by a quantum gravitational bounce. In classical Einstein theory, the angular tidal forces close to shell-crossing singularities do not diverge, while despite radial tidal forces being divergent, the radial deviation does not go to zero at SCS, nor does it diverge Szekeres:1995gy . This makes such singularities less pathological than the central one. Here, we show that similar results also hold at the effective level.
To prove this, we compute explicitly (16), (17), using the solution (3) of the effective equation (2). Let us start from the equation for the angular deviation. After some computation, we find
@@
Firstly, we note that in the classical case $\Delta\rightarrow 0$, the angular tidal forces negatively diverge for $t\rightarrow\alpha(R)$ (classical central singularity). In contrast, here it remains always bounded, reaches its maximum at $t=\alpha(R)$ (the bouncing point) and has two zeros, for $t=\alpha\pm 2\sqrt{\Delta}/\sqrt{3}$.
@@
Figure 1: Angular tidal force for $\alpha=1$, in the classical (dark blue curve) and effective collapse (red curve). The dotted vertical line signals the latest time at which a SCS occurs in the effective case.
In the classical case the function diverges at the crush singularity $t=\alpha=1$, while remains bounded at the effective level.
Fig.1 shows the comparison between the classical and effective cases. Notice that angular tidal forces do not diverge at shell-crossing singularity. Although the latest time at which a SCS arises and the largest zero of (18) are quantitatively similar, they must be considered as two distinct features.
@@
Let us now examine the radial tidal force. By explicitly writing the right-hand side of equation (16) (divided by $\delta x^{1}$), after a lengthy but straightforward computation, we find (see Appendix A)
@@
Firstly, we notice that divergence can only occur if $\partial_{R}r=0$, provided that the numerator of the last term is non-zero. This indicates that the radial tidal force remains finite at the bounce, while it instead diverges at SCS. To study the radial tidal force near and at SCS, we can assume that the first term is negligible
@@
Next, we recall that for the profiles we are considering (decreasing, continuous, and of compact support), shell-crossing singularities arise for $\alpha<t\leq\alpha+2\sqrt{\Delta}/3$, which implies $F\Delta<r^{3}\leq 2F\Delta$, from (3). This means that the numerator of (21) is strictly positive and that at the shell-crossing singularity the radial tidal force diverges. It is worth noting that such a divergence arises in the deep quantum gravitational regime and is a positive divergence: an infinite stretching radial force arises, conceptually similar to the one occurring classically near the central singularity, producing the spaghettification effect. In contrast, near classical shell-crossing singularities, the radial tidal force is negatively divergent: two points radially separated tend to be crushed together by the tidal force, even though in a weak fashion Szekeres:1995gy , as one can see by setting $\Delta=0$ in (21).
Let us now focus on the (positive) divergent behavior of the radial tidal force near and at the SCS. The condition for SCS can be derived from the solution (3) and rewritten as Fazzini:2023ova
$(t-\alpha)^{2}\partial_{R}m-2m(t-\alpha)\partial_{R}\alpha+\frac{4\Delta}{9}\partial_{R}m=0~,$ (22)
see also Appendix B. This is a second-degree equation in $t-\alpha$, with solutions $t_{1}=\alpha(R)+\delta t_{1}(R)$, $t_{2}=\alpha(R)+\delta t_{2}(R)$, with $\delta t_{1}\leq\delta t_{2}$. These solutions are guaranteed to be real and positive for continuous, inhomogeneous, decreasing profiles (for which $\partial_{R}\alpha>0$ Fazzini:2025hsf ) of compact support, as a consequence of the results in the previous section. It can also be easily shown that the solutions are distinct Fazzini:2023ova , except for an initial condition that satisfies $\frac{m\partial_{R}\alpha}{\partial_{R}m}=\frac{2\sqrt{\Delta}}{3}$, which is a fine-tuned case which we do not consider here. Therefore, we can rewrite the condition for SCS as
@@
This means that close to $t=t_{1}=\alpha+\delta t_{1}$, the function $r^{\prime}$ goes to zero linearly: $\partial_{R}r\sim K(t_{1},t_{2},R)(t_{1}-t)$ for $t\lesssim t_{1}$, with $K(t_{1},t_{2},R)>0$ (see Appendix B). Here we are only interested in the tidal forces before and at the shell-crossing singularity. Beyond this point, the analytic solution (3) is not reliable. With this in hand, we can rewrite (16):
$\partial^{2}_{t}\delta x^{1}\sim\frac{B(t_{1},t_{2},R)}{t_{1}-t}\delta x^{1}~,$ (24)
where $B(t_{1},t_{2},R)$ is given by $\partial^{2}_{t}\partial_{R}r$ evaluated at shell-crossing singularity divided by $K(t_{1},t_{2},R)$. It is important to keep in mind that this approximation holds only close to the SCS ($t\lesssim\alpha(R)+\delta t_{1}(R)$). Equation (24) can be solved analytically. Let us call $t_{1}=t_{SCS}$, and define $\delta t\equiv t_{SCS}-t>0$. The solution is
$\delta x^{1}(\delta{t})\sim$ $c_{1}(R)\sqrt{\delta{t}}~I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]+$ (25)
$c_{2}(R)\sqrt{\delta{t}}~K_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]~,$ (26)
where $I_{1}$ and $K_{1}$ are modified Bessel functions of the first and second kinds, and $c_{1}(R),c_{2}(R)$ are time independent functions determined by initial conditions. The previous solution holds only for $\delta{t}\rightarrow 0^{+}$. While the first term goes to zero in this limit, the second term approaches a constant non-zero value. Their explicit Taylor expansions are:
@@
Therefore, even though the radial tidal forces diverge at SCS, two neighboring points that are radially separated (and thus belong to two different matter layers) do not generally clash when a shell-crossing singularity occurs. For example, these points can be thought of as belonging to two shells close to the ones that cross, but with a larger coordinate separation $\delta R$). If these points belong exactly to the crossing shells, they will be crushed, and this condition is given by $c_{2}(R)=0$ Szekeres:1995gy , since at SCS $\partial_{R}r(R,t_{SCS})=0$, that implies $\delta r=\delta x^{1}=0$. This implies that the radial portions of the collapsing star (or even geodesic test particles radially separated) undergo spaghettification, but without being infinitely stretched. We refer to this effect as gentle spaghettification.
This result leads to the conclusion that SCSs, which generally arise in effective stellar collapse, are less pathological than classical crush singularities. It also justifies the extension of the spacetime beyond SCS Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf ; Liu:2025fil ; Fazzini:2025zrq ; Husain:2025wrh . It is important to emphasize that this result is nontrivial. As mentioned above, even though the structure of the deviation equation follows the classical form, the effective solution differs from the classical one. Therefore, the behavior of tidal forces at shell-crossing singularities also differs from the classical behavior in the effective model.
To conclude, for profiles that develop shell-crossing singularities in the pre-bounce phase (where the classical Einstein equations hold), the classical analysis Szekeres:1995gy can be directly exported to the effective model.
@@
Shell-crossing singularities are physical singularities generally unavoidable in effective dust collapse inspired by loop quantum gravity, within the $\bar{\mu}+K$-loop quantization scheme. They therefore play a central role in this effective model, and their geometrical and physical features need to be investigated. In this work, we focus on the analysis of the tidal forces arising from effective dust collapse, with particular attention to the bouncing point and shell-crossing singularities. We found that tidal forces remain bounded at the bounce, in contrast to what happens in the classical case at the central crush singularity, and the same holds for the angular component at the shell-crossing singularity. However, the radial component of the tidal force positively diverges at the shell-crossing singularity.
@@
Nevertheless, when studying the behavior of the radial deviation by solving the Jacobi equation, we find that it remains bounded throughout the dynamics, also when the shell-crossing singularity arises. The positive divergent behavior of the tidal force implies a spaghettification effect, similar to the one occurring near the central singularity in classical stellar collapse. However, the bounded behavior of the radial displacement between nearby points makes such an effect much less dramatic. We refer to this phenomenon as gentle spaghettification, and it allows the dynamics to extend beyond shell-crossing singularities in a meaningful way, through the integral form of the equations of motion. Future work in this direction will aim to investigate the tidal forces in the effective gravitational collapse of dust in the non-marginally bound case Cipriani:2024nhx , and of fluids with pressure Cafaro:2024lre , in order to understand, at least at the numerical level, the role of pressure in the tidal forces experienced by test particles.
Although this work does not aim to study the spacetime evolution beyond the formation of shell-crossing singularities, the analysis presented here could be applied to spacetimes where such an extension has been performed, particularly to investigate tidal forces near the non-isolated thin shell generated by the SCS. When SCSs develop, the effective equations expressed in Painlevé-Gullstrand (PG) coordinates exhibit characteristic crossing, which imply a multi-valued behavior of the extrinsic curvature. Since these equations take the form of hyperbolic conservation laws in the marginally bound case Husain:2022gwp and balance laws in the non-marginally bound case Cipriani:2024nhx , one might be tempted to study their dynamics through weak solutions, as commonly done in other fields where characteristic crossings occur Batchelor ; Hoefer2006 . This approach results in an evolving shock in the extrinsic curvature Husain:2022gwp ; Cipriani:2024nhx .
However, it was recently shown Fazzini:2025zrq that the integral formulation underlying weak solutions implicitly assumes the continuity of PG time throughout the shock (the SCS in LTB coordinates), which is physically questionable. Indeed, the Misner-Sharp mass function becomes discontinuous at the shock, implying different time dilations measured by interior and exterior metrics, and consequently different PG times flowing. Furthermore, weak solutions exhibit superluminal behavior. The proper way to extend the spacetime beyond the SCS is to apply the Israel junction conditions to a non-isolated thin shell, ensuring subluminal propagation and allowing PG time to be discontinuous at the shock. Once this framework is established, understanding the evolution of tidal forces near the evolving thin shell will be crucial.
To conclude, a natural question arises as to whether shell-crossing singularities can be avoided once a full quantum theory of dust collapse is taken into account, going beyond the effective approach. In principle, quantum gravitational fluctuations in the fields could introduce corrections to the effective equation (2) that prevent the formation of SCS and lead to a singularity-free evolution, although this outcome is not guaranteed. If this scenario holds, the thin shell dynamics beyond SCS studied here might qualitatively capture the behavior predicted by a more refined effective model. Moreover, since quantum fluctuations are expected to decay rapidly near the SCS, the tidal forces computed in this work could provide a reasonable approximation to those in such a hypothetical effective model.
@@ Acknowledgements.
Appendix A The radial tidal force
We can compute the quantity $\frac{\partial_{t}^{2}\partial_{R}r}{\partial_{R}r}$ from the general solution (3). By deriving twice this expression in time, and rewriting the result in term of $r(R,t)$, we get
$\partial_{t}^{2}r(R,t)=\frac{F(R)^{\frac{2}{3}}}{r^{4}}\left[r^{3}+2\Delta F(R)\right]~.$ (29)
Then, deriving with respect of the radial coordinate, one gets
$\partial_{R}{\partial_{t}^{2}r(R,t)}=\frac{2\Delta}{r^{4}}\partial_{R}F^{\frac{5}{3}}+\frac{\partial_{R}F^{\frac{2}{3}}}{r}-F^{\frac{2}{3}}\left(1+\frac{8\Delta F}{r^{3}}\right)\frac{\partial_{R}r}{r^{2}}~.$ (30)
By dividing this expression by $\partial_{R}r(R,t)$, and rearranging terms, we get
Appendix B Behavior of $\partial_{R}r(R,t)$ close to SCS
In this appendix, we prove that, close to SCS $t\lesssim t_{1}$, we can write $\partial_{R}r(R,t)=K(t_{1},t_{2},R)(t_{1}-t)+O(t_{1}-t)$, where $t_{1}(R)$ and $t_{2}(R)$ are the two real roots of (22), and $K(t_{1},t_{2},R)$ is a positive function. We can start by differentiating (3) with respect to $R$, getting
$\frac{\partial_{R}r}{r}=\frac{\partial_{R}m}{3m}+\frac{3}{2}\frac{(\alpha(R)-t)\partial_{R}\alpha}{\left[\frac{9}{4}(\alpha-t)^{2}+\Delta\right]}~,$ (32)
which can be rewritten as
$\partial_{r}(R,t)=\frac{3r}{4m}\frac{\partial_{R}m(\alpha-t)^{2}+\frac{4}{9}\Delta\partial_{R}m-2\partial_{R}\alpha(t-\alpha)}{\frac{9}{4}(t-\alpha)^{2}+\Delta}~.$ (33)
We recognize the last term in the numerator as the left-hand side of (22), and therefore we can rewrite the previous equation as
$\partial_{R}r(R,t)=\frac{3r}{4m}\frac{[t-t_{1}(R)][t-t_{2}(R)]}{\frac{9}{4}(t-\alpha)^{2}+\Delta}~,$ (34)
where $t_{1}(R)$ and $t_{2}(R)$ are the solutions of (22). Now, let us study this expression for $t-t_{1}=-\varepsilon$, $\varepsilon>0$ (close but before SCS). By explicitly writing $r(R,t)$ using (3), we get
$\partial_{R}r(R,t_{1}-\varepsilon)=\frac{3F^{\frac{1}{3}}}{4m(R)}\frac{\varepsilon(\varepsilon+t_{2}(R)-t_{1}(R))}{\left[\frac{9}{4}(t_{1}-\alpha(R)-\varepsilon)^{2}+\Delta\right]^{\frac{2}{3}}}~,$ (35)
and, for $\varepsilon<<1$:
$\partial_{R}r(R,t_{1}-\varepsilon)=\frac{3F^{\frac{1}{3}}}{4m(R)}\frac{\varepsilon(t_{2}(R)-t_{1}(R))}{\left[\frac{9}{4}(t_{1}-\alpha(R))^{2}+\Delta\right]^{\frac{2}{3}}}+O(\varepsilon)~.$ (36)
If we define
$K(t_{1},t_{2},R)\equiv\frac{3F}{4m}\frac{(t_{2}(R)-t_{1}(R))}{\left[\frac{9}{4}(t_{1}-\alpha(R))^{2}+\Delta\right]^{\frac{2}{3}}}$ (37)
then we can write
$\partial_{R}r(R,t)=K(t_{1},t_{2},R)(t_{1}-t)+O(t_{1}-t)~,$ (38)
for $t\lesssim t_{1}$, with $K(t_{1},t_{2},R)>0$.